OpenStudy (anonymous):

please help... If 5x+4y=21 and x+y=9, then 6x+5y=

OpenStudy (anonymous):

You have to use the first two equations to find x and y, then use this answer to solve the last equation. \[5x+4y=21 ...(1)\]\[x+y=9...(2)\]From (2), we have y = 9 - x. Substitute this into (1) to obtain\[5x+4(9-x)=21 \rightarrow 5x+36-4x=21 \rightarrow x = -15\]If x = -15, we can use the result from (2) to obtain \[y=9-x=9-(-15)= 24\] Hence,\[6x+5y=6(-15)+5(24)=30\]

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

^^ Same here :)